# Roman Garipov ${}^{*\,\dagger}$

### v1

@@ Abstract
We introduce AutoJudge [^1] Our code is available at github.com/garipovroma/autojudge., a framework that accelerates large language model (LLM) inference with task-specific lossy speculative decoding. Instead of matching the original model output distribution token-by-token, we identify which of the generated tokens affect the downstream quality of the generated response, relaxing the guarantee so that the “unimportant” tokens can be generated faster. Our approach relies on a semi-greedy search algorithm to test which of the mismatches between target and draft models should be corrected to preserve quality, and which ones may be skipped. We then train a lightweight classifier based on existing LLM embeddings to predict, at inference time, which mismatching tokens can be safely accepted without compromising the final answer quality. We test our approach with Llama 3.2-1B (draft) and Llama 3.1-8B (target) models on zero-shot GSM8K reasoning, where it achieves up to 1.5×1.5$\times$1.5 × more accepted tokens per verification cycle with under 1% degradation in answer accuracy compared to standard speculative decoding and over 2×2$\times$2 × with a small loss in accuracy. When applied to the LiveCodeBench benchmark, our approach automatically detects other, programming-specific important tokens and shows similar speedups, demonstrating its ability to generalize across tasks.
@@ 1 Introduction
Recent advances in LLM capabilities, including chain-of-thought reasoning (Wei et al., 2022; Kojima et al., 2022; Suzgun et al., 2022), writing complex software (Rozière et al., 2023; Li et al., 2023; Jiang et al., 2024), or interacting with external tools (Schick et al., 2023; Qin et al., 2023), increasingly rely on inference-time computation (Snell et al., 2024; Beeching et al., 2024). This progress is further accelerated with the release of reasoning-capable models, both proprietary (OpenAI et al., 2024; Anthropic, 2024; Google DeepMind, 2025) and open-access (DeepSeek-AI et al., 2025; Meta, 2025; Qwen Team, 2025), that were explicitly trained to perform these kinds of inference-time computation. However, as the LLMs tackle harder problems, they also tend to generate longer sequences (Muennighoff et al., 2025) with tens of thousands of tokens (Yeo et al., 2025), taking up tens of minutes (and hundreds of dollars) per task (ARC Prize Foundation, 2024).
A popular way to speed up LLM inference is through speculative decoding (Leviathan et al., 2023; Chen et al., 2023) that uses a small “draft” model to propose the likely next tokens, then verifies these tokens with the main model in parallel. Speculative decoding and its successors (Miao et al., 2023; Cai et al., 2024; Li et al., 2024b) can speed up LLM inference while guaranteeing that the generated outputs match the original model (for greedy inference) or follow the same sampling distribution. To achieve this, speculative decoding algorithms check if the draft tokens match the original model predictions. If there is a mismatch, they discard the incorrect token and all subsequent ones.
@@
In this work, we look for ways to streamline this process. Instead of relying on human annotators, we propose AutoJudge: a search-based algorithm that detects which tokens are important for the task at hand based on how they affect the final answer. The algorithm is based on the idea that a token cannot be deemed “important” by itself, but in combination with other generated tokens. Thus, we propose a procedure that selects a small subset of important mismatching tokens that affect the final answer. Using this procedure, we can automatically mine a dataset to train an important token classifier that can then be used to accelerate speculative decoding.
We evaluate our approach on two problem types: mathematical reasoning and programming. In each case, the proposed search algorithm finds a small set of task-specific contextual “important tokens” — situations where main and draft models disagree on the next token in a way that affects the final response quality. We then train a classifier to detect these important tokens and use it to improve traditional speculative decoding by relaxing its verification procedure. Our experiments with Llama 3.x models demonstrate that the proposed approach can accept over 15 tokens per target model forward pass (up to 1.5×1.5$\times$1.5 × that of speculative decoding) at the cost ${\leq}1{\%}$ drop in accuracy on GSM8K (Cobbe et al., 2021) and over 20 tokens with minor accuracy drawdown. When applied to programming tasks on LiveCodeBench (Jain et al., 2024), our approach is able to determine different task-specific important tokens, showing similar performance gains. The proposed framework is simple and general, using a classifier only when the original algorithm would reject a token, making it compatible with arbitrary speculative decoding algorithms.
Figure 1: Intuitive scheme of the proposed approach: (left) data gathering: detecting mismatching tokens that affect final response quality; these tokens are then used to train a classifier (right) using the trained classifier to generate more tokens per cycle with speculative decoding.
@@ 2 Background
Our work builds on top of speculative decoding (Stern et al., 2018; Leviathan et al., 2023; Chen et al., 2023), a family of inference algorithms that accelerate token generation by improving hardware utilization. Speculative Decoding uses an auxiliary “draft” model to generate $K{>}1$ possible future tokens, then runs the main “target” model in parallel to verify [^1] For greedy decoding, it checks that the drafted tokens are the same as the target model’s own next token predictions. For sampling, it uses a procedure that matches sampling probabilities (Leviathan et al., 2023). the generated tokens. The drafted tokens that agree with the target model predictions are accepted by the algorithm. In turn, the first mismatching token and all subsequent ones are rejected. This way, the method guarantees that all generated tokens follow the same distribution as sampling from the target model.
Subsequent works improve on this idea by generating draft trees instead of single sequences (Miao et al., 2023; Liu et al., 2023; Chen et al., 2024; Svirschevski et al., 2024), training specialized “heads” to draft next tokens based on the model’s hidden states (Cai et al., 2024; Ankner et al., 2024; Li et al., 2024b, a), and more (Fu et al., 2023; Spector and Re, 2023; Sun et al., 2023; He et al., 2023).
@@
The main challenge of Judge Decoding is determining which of the generated tokens can be skipped this way. Bachmann et al. (2025) address this problem by manually labeling a training dataset for the classifier. Judge Decoding requires human annotators to find the “mistake” — the first mismatching token that led the draft model to diverge from the original answer. The resulting dataset of high-quality training examples is then used to train a linear classifier that detects such “mistakes” during inference.
Authors demonstrate that the collected dataset can, in principle, be reused for different tasks and models. However, using the dataset gathered from one task for inference on a different task results in substantial performance drawdown. Intuitively, different tasks (such as creative writing, math, or programming) have different criteria for which parts of the generated response matter most. Hence, it is best to train the important token classifier for the exact task at hand. However, doing so with Judge Decoding would require relabeling the dataset by human annotators, which can be costly and time-consuming if the task domain requires specialized expertise such as medicine or law. To alleviate this problem, we develop an automated search procedure for determining important tokens without external human (or LLM) annotators.
3 Method
@@ 3.1 Mining Important Tokens
3:  $\mathcal{M}\leftarrow\emptyset$                             $\triangleright$ A set of tuples (position, target token, draft token, important)
4:  $y\leftarrow\textsc{generate}(x,\theta_{\text{target}})$
5:  $\alpha\leftarrow\textsc{extractAnswer}(y)$
@@
7:  $\mathcal{I}\leftarrow\{i\mid y_{i}\neq\widetilde{y}_{i}\}$        $\triangleright$ Indices where draft and target tokens mismatch
@@
10:     $\hat{y}=y_{1:t}\oplus\widetilde{y}_{t}\oplus\textsc{generate}(x\oplus y_{1:t}\oplus\widetilde{y}_{t},\theta_{target})$triangleright Replace $\widetilde{y}_{t}$ and continue with $\theta_{target}$
11:     $\hat{\alpha}\leftarrow\textsc{extractAnswer}(\hat{y})$
@@
13:        $\mathcal{M}\leftarrow\mathcal{M}\cup\{(t,y_{t},\widetilde{y}_{t},\texttt{False})\}$triangleright Equivalent answer, token $y_{t}$ is not important
@@
[…]
@@
17:        $\mathcal{M}\leftarrow\mathcal{M}\cup\{(t,y_{t},\widetilde{y}_{t},\texttt{True})\}$triangleright Different answer, token $y_{t}$ is important, keep it
@@
[…]
Following this notation, let $y_{target}=\textsc{generate}(x,\theta_{target})$ be the main model outputs. A token $y_{t}\in y_{target}$ is unimportant if swapping that token for the draft model output results in an equivalent answer. Likewise, if replacing $y_{i}$ (and continuing target generation from there) results in a different answer, then the original token was “important” and the token should be generated with $\theta_{target}$.
@@
where $y_{1:t}=y_{1},\dots,y_{t-1}$ denotes taking a prefix of $y$ up to, but excluding index $t$.
[…]
When deciding if a mismatching token is important for the final response, we need to account for the fact that changing one token will most likely lead to changes in subsequent tokens. A naïve way to account for that change is by continuing [^2] To simplify notation, we assume that the $\textsc{generate}(\cdot,\cdot)$ function can be called with a prefix of a response. In that case, we assume that the total response length (and not just newly generated tokens) does not exceed $T_{max}$, so that the response cannot grow indefinitely with each subsequent replacement. the response after replacing one token $\widetilde{y}_{t}$:
@@
However, this approach has a significant downside in that it assumes that all subsequent tokens will be generated by $\theta_{target}$, whereas in reality, some of them may be generated by $\theta_{draft}$ following the same algorithm. In preliminary experiments, we found that, with a capable-enough $\theta_{target}$, even significant generation mistakes can be detected and self-corrected (similar to the “Aha moment” from DeepSeek-AI et al. (2025); Muennighoff et al. (2025)). However, if the model makes multiple mistakes, they eventually reach a critical mass, leading to an incorrect answer.
To address this, we re-frame our task from detecting individual important tokens to finding combinations of tokens that jointly affect the final answer. This changes our problem to finding the minimal set of mismatching tokens that need to be generated by $\theta_{target}$ while still producing an equivalent answer [^3] More precisely, find the fastest-to-generate sequence that produces an equivalent answer, accounting for the differences in response length.. Since replacing a single mismatching token affects all subsequent token choices, the exact solution to this problem requires a tree search over possible token assignments. While type of tree search is possible, it would take up significant runtime due to the large number of LLM forward passes required to try all mismatch combinations.
To simplify the procedure, we opt instead for a simpler, semi-greedy search that starts from the target model response and iteratively tries to replace mismatching target model tokens with their draft counterparts. If replacing a token affects the final answer, we deep this token important and keep the original (target model) version. If, however, replacing the token results in an equivalent answer, we deem this token unimportant, replace it with the draft model version and continue the search from the new sequence, with a different suffix and possibly a different $\mathcal{I}$. That way, we guarantee that the search algorithm is aligned with what happens during inference: the important tokens are generated with the target model and the unimportant ones are kept from the draft model. We summarize the resulting search procedure in Algorithm 1 and discuss some of its implications in Appendix A.
@@
To find the total number of animal legs, we need to calculate the legs$\,{}_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}\texttt{[total]}}}$ of each animal and then add them up.
- 2 dogs have 4$\,\,{}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}[2]}}$ legs each, so 2 dogs have 2 *$\,{}_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}\texttt{[times]}}}$ 4 = 8 legs.
@@
Now$\,{}_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}\texttt{[Adding]}}}$, let’s add the legs together$\,{}_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}\texttt{[of]}}}$ , we get 8 (from the dogs) + 2 (from the chicken) = 10 legs.
@@
# -*-${}_{[\texttt{YOUR}]}$ coding: utf-8 -*-
${}_{[\texttt{\#}]}$ solve${}_{[\texttt{check}]}$(s):
@@
3.2 Classifier Training
Once we gather a dataset of task-specific important tokens with Alg. 1, we can train a classifier that would detect such tokens for use during inference. This classifier can, in principle, be any type of model, from a simple linear model or decision tree to a fine-tuned transformer layer. However, in our work, we default to training lightweight linear models with existing LLM hidden states as features, since those would introduce the least overhead during inference. There are several important design choices that can affect the effectiveness of such classifier: we address each one separately.
1. Which token representations to use? The hidden states that predicted the mismatched token, or the next hidden states that encode the mismatched token itself? In our experiments, we found that using the latter representations results in substantially greater classifier accuracy (see Appendix B). However, obtaining these representations comes with a caveat.
@@ 3.2 Classifier Training
2. Which token alternative to use? Since the classifier works best with the representations from encoding the mismatching token, it is natural to ask which token should be encoded: the draft token, the mismatching target token, or both? When analyzing this, we found that using both token representations comes with slight increase in classifier accuracy (see Appendix B). However, obtaining these representations in practice would require running $\theta_{target}$ more than once during the verification stage, which would complicate inference and introduce performance overhead. For this reason, we opt to only use use the draft token representations for the classifier, since those are already available in normal speculative decoding.
3. Which model provides feature representations? During verification stage, we have access to both draft and target model representations: we can use either or both of them as inputs. In practice, we found that using both draft and target model representations (concatenated) gives slightly better results than target model, and using draft model representations alone is substantially worse. Since both representations are already available during inference, we opt to use both representations.
Classifier model & training. In this work, we train a simple logistic regression to detect important tokens. While a more complex model could achieve greater accuracy, logistic regression is significantly easier to deploy, has less runtime & memory overhead and needs less training data. Furthermore, it can be fused with the existing “LM head” layer of the draft and target LLMs, which would make its computation virtually free. To control overfitting, we perform a simple grid search over the $L_{2}$ regularization coefficient (“$C$”) with a logarithmic grid. We report additional details in Appendix B.

### v2

@@ Abstract
We introduce AutoJudge [^1] Our code is available at github.com/garipovroma/autojudge., a method that accelerates large language model (LLM) inference with task-specific lossy speculative decoding. Instead of matching the original model output distribution token-by-token, we identify which of the generated tokens affect the downstream quality of the response, relaxing the distribution match guarantee so that the “unimportant” tokens can be generated faster. Our approach relies on a semi‑greedy search algorithm to test which of the mismatches between target and draft models should be corrected to preserve quality and which ones may be skipped. We then train a lightweight classifier based on existing LLM embeddings to predict, at inference time, which mismatching tokens can be safely accepted without compromising the final answer quality. We evaluate the effectiveness of AutoJudge with multiple draft/target model pairs on mathematical reasoning and programming benchmarks, achieving significant speedups at the cost of a minor accuracy reduction. Notably, on GSM8k with the Llama 3.1 70B target model, our approach achieves up to ${\approx}2{\times}$ speedup over speculative decoding at the cost of ${\leq}1\%$ drop in accuracy. When applied to the LiveCodeBench benchmark, AutoJudge automatically detects programming-specific important tokens, accepting ${\geq}25$ tokens per speculation cycle at $2\%$ drop in Pass@1. Our approach requires no human annotation and is easy to integrate with modern LLM inference frameworks.
@@ 1 Introduction
Recent advances in LLM capabilities, including chain-of-thought reasoning (Wei et al., 2022; Kojima et al., 2022; Suzgun et al., 2022), writing complex software (Rozière et al., 2023; Li et al., 2023; Jiang et al., 2024), or interacting with external tools (Schick et al., 2023; Qin et al., 2023), increasingly rely on inference-time computation (Snell et al., 2024; Beeching et al., 2024). This progress is further accelerated with the release of reasoning-capable models, both proprietary (OpenAI et al., 2024; Anthropic, 2024; Google DeepMind, 2025) and open-access (DeepSeek-AI et al., 2025; Meta, 2025; Qwen Team, 2025), that were explicitly trained to perform these kinds of inference-time computation. However, as LLMs tackle harder problems, they also tend to generate longer sequences (Muennighoff et al., 2025) with tens of thousands of tokens (Yeo et al., 2025), taking up tens of minutes (and hundreds of dollars in costs) per task (ARC Prize Foundation, 2024).
A popular way to speed up LLM inference is through speculative decoding (Leviathan et al., 2023; Chen et al., 2023), which uses a small “draft” model to propose the likely next tokens, then verifies these tokens with the main model in parallel. This method, along with its successors (Miao et al., 2023; Cai et al., 2024; Li et al., 2024b), can speed up LLM inference while guaranteeing that the generated outputs match those of the original model (for greedy inference) or follow the same distribution. To achieve this, speculative decoding algorithms check if the draft tokens match the original model predictions. If there is a mismatch, they discard the incorrect token and all subsequent ones.
@@
In this work, we look for ways to streamline this process. Instead of relying on human annotators, we propose AutoJudge, a search-based algorithm that detects which tokens are important for the task at hand based on how they affect the final answer. The algorithm is based on the idea that a token is deemed “important” not by itself, but in combination with other generated tokens. Thus, we propose a procedure that selects a small subset of important mismatching tokens that affect the final answer. Using this procedure, we can automatically mine a dataset to train an important token classifier that can then be used to accelerate speculative decoding.
The proposed search algorithm finds a small set of task-specific contextual “important tokens” — cases where target and draft models disagree on the next token in a way that affects the final response quality. We then train a classifier to detect these important tokens and use it to improve traditional speculative decoding by relaxing its verification procedure.
Our experiments with Llama 3.x models demonstrate that the proposed approach can indeed identify important tokens and save time on speculation, accepting on average over 40 tokens per target model forward pass, (approx. $2{\times}$ that of speculative decoding) at the cost ${\leq}1{\%}$ drop in accuracy on GSM8K (Cobbe et al., 2021) and even more with minor accuracy drawdown. We obtain similar results with the Qwen2.5(Yang et al., 2024) family models, observing comparable number of accepted tokens and accuracy trade-offs. When applied to programming tasks on LiveCodeBench (Jain et al., 2024), our approach is able to determine different task-specific important tokens, showing similar performance gains. The proposed framework is simple and general, using a classifier only when the original algorithm would reject a token, making it compatible with arbitrary speculative decoding algorithms. The main contributions of our work can be summarized as follows:
  • We formulate AutoJudge, an algorithm for detecting which of the tokens generated in speculative decoding affect the downstream quality of a given task. Our algorithm requires no human annotation and can be applied to most popular LLM tasks.
  • We verify the efficacy of AutoJudge on mathematical reasoning and programming benchmarks for several speculative decoding setups (model pairs) for both mathematical reasoning and programming. Our evaluations demonstrate favorable tradeoffs between accuracy and inference speedup, generating 20-45 tokens per speculative decoding cycle at the cost of slight accuracy drawdown.
  • We integrate our inference procedure with the vLLM library (Kwon et al., 2023) and report inference speed on A100 GPUs for both 8B and 70B target models, with up to $2{\times}$ speedup over speculative decoding at ${\leq}1\%$ quality drawdown, and on H100 GPUs with a 405B target model.
Figure 1: Intuitive scheme of the proposed approach: (left) data collection: detecting mismatching tokens that affect final response quality; these tokens are then used to train a classifier (right) using the trained classifier to generate more tokens per cycle with speculative decoding.
@@ 2 Background
Our work builds on top of speculative decoding (Stern et al., 2018; Leviathan et al., 2023; Chen et al., 2023), a family of inference algorithms that accelerate token generation by improving hardware utilization. Speculative Decoding uses an auxiliary “draft” model to generate K>1K{>}1 possible future tokens, then runs the main “target” model in parallel to verify∗*∗**For greedy decoding, it checks that the drafted tokens are the same as the target model’s own next token predictions. For sampling, it uses a procedure that matches the sampling probabilities (Leviathan et al., 2023). the generated tokens. The drafted tokens that agree with the target model predictions are accepted by the algorithm. In turn, the first mismatching token and all subsequent ones are rejected. This way, the method guarantees that all generated tokens follow the same distribution as sampling from the target model. Subsequent works improve on this idea by generating draft trees instead of single sequences (Miao et al., 2023; Liu et al., 2023; Chen et al., 2024; Svirschevski et al., 2024), training specialized “heads” to draft next tokens based on the model’s hidden states (Cai et al., 2024; Ankner et al., 2024; Li et al., 2024b, a), and more (Fu et al., 2023; Spector and Re, 2023; Sun et al., 2023; He et al., 2023).
@@
The main challenge of Judge Decoding is determining which of the generated tokens can be skipped this way. Bachmann et al. (2025) address this problem by manually labeling a training dataset for the classifier. Judge Decoding requires human annotators to find the “mistake” — the first mismatching token that led the draft model to diverge from the original answer. The resulting dataset of high-quality training examples is then used to train a linear classifier that detects such “mistakes” during inference. Authors demonstrate that the collected dataset can, in principle, be reused across tasks and models. However, using the data from one task for inference on a different task results in substantial performance drawdown. Intuitively, different tasks (such as creative writing, math, or programming) have different criteria for which parts of the generated response matter most. Hence, it is best to train the important token classifier for the exact task at hand. However, doing so with Judge Decoding would require relabeling the data by human annotators, which can be costly and time-consuming in specialized domains such as medicine or law. To alleviate this problem, we develop an automated search procedure for determining important tokens without external human (or LLM) annotators.
3 Method Overview
@@ 3.1 Mining Important Tokens
[…]
@@
10:  $\hat{y}=y_{1:t}\oplus\widetilde{y}_{t}\oplus\textsc{generate}(x\oplus y_{1:t}\oplus\widetilde{y}_{t},\theta_{target})$ $\triangleright$ Replace $\widetilde{y}_{t}$ and continue with $\theta_{target}$
[…]
@@
13:   $\mathcal{M}\leftarrow\mathcal{M}\cup\{(t,y_{t},\widetilde{y}_{t},\texttt{False})\}$ $\triangleright$ Equivalent answer, token $y_{t}$ is not important
@@
15:   $\widetilde{y}\leftarrow\textsc{forward}(x\oplus y,\theta_{\text{draft}})\texttt{.argmax(-1)[len(x)-1:-1]}$
@@
17:   $\mathcal{M}\leftarrow\mathcal{M}\cup\{(t,y_{t},\widetilde{y}_{t},\texttt{True})\}$ $\triangleright$ Different answer, token $y_{t}$ is important, keep it
@@
Without the loss of generality, we also assume that there is a problem-specific way to extract the final answer from the model’s response, $a=\textsc{extractAnswer}(y)$. In mathematical reasoning tasks such as GSM8K Cobbe et al. (2021), the final answer is literally whatever the model puts after "the final answer (is)". In programming tasks, the “answer” would be the output from the testing system given the generated code — either a report about passing and failing tests or a testing error (e.g., an Out Of Memory or Syntax Error). Finally, we say that two answers are equivalent $a_{ref}{\equiv}a_{alt}$ if they are the same from the downstream task perspective. Note that this does not require them to be exactly equal as strings: in math problems, $\texttt{1.5}{\equiv}\texttt{3/2}$, whereas in programming tasks, two programs can be equivalent despite having different variable names. If the task at hand does not have a formalized evaluation procedure, e.g., general conversation agents, we can define $\textsc{extractAnswer}(y)=y$ and detect if two answers are equivalent using an LLM or human judges.
Following this notation, let $y_{target}=\textsc{generate}(x,\theta_{target})$ be the main model outputs. A token $y_{t}\in y_{target}$ is unimportant if swapping that token for the draft model’s output results in an equivalent answer. Likewise, if replacing $y_{i}$ (and continuing target generation from there) results in a different answer, then the original token was “important” and the token should be generated with $\theta_{target}$.
@@
where $y_{1:t}=y_{1},\dots,y_{t-1}$ denotes taking a prefix of $y$ up to, but excluding index $t$. In practice, we can find these tokens quickly by re-encoding the target model response with the draft model: $\textsc{forward}(x\oplus y,\theta_{draft}).\texttt{argmax(dim=-1)[M{-}1:M{+}T{-}1]}$ , where $x\oplus y$ denotes concatenation, $\textsc{forward}(\cdot,\cdot)$ is a parallel transformer forward pass that outputs next token logits, and the $\texttt{logits}.\texttt{argmax(dim=-1)[M{-}1:M{+}T{-}1]}$ takes the most likely next tokens for every position, excluding the prompt and accounting for the shift from next token prediction.
When deciding if a mismatching token is important for the final response, we need to account for the fact that changing one token will most likely lead to changes in subsequent tokens. A naïve way to account for that change is by continuing∗*∗**For notation simplicity, we assume that the $\textsc{generate}(\cdot,\cdot)$ function can be called with a prefix of a response. In that case, we assume that the total response length (and not just newly generated tokens) does not exceed $T_{max}$, so that the response cannot grow indefinitely with each subsequent replacement. the response after replacing one token $\widetilde{y}_{t}$:
@@
However, this approach has a significant downside in that it assumes that all subsequent tokens will be generated by $\theta_{target}$, whereas in reality, some of them may be generated by $\theta_{draft}$ following the same algorithm. In preliminary experiments, we found that, with a capable enough $\theta_{target}$, even significant generation errors can be detected and self-corrected (similar to the “Aha moment” from DeepSeek-AI et al. (2025); Muennighoff et al. (2025)). However, if the model makes multiple mistakes, they eventually reach a critical mass, leading to an incorrect answer.
To address this, we reframe our task from detecting individual important tokens to finding combinations of tokens that jointly affect the final answer. This changes our problem to finding the minimal set of mismatching tokens that need to be generated by $\theta_{target}$ while still producing an equivalent answer∗*∗**More precisely, find the fastest-to-generate sequence, accounting for the differences in response length.. Since replacing a single mismatching token affects all subsequent token choices, the exact solution to this problem requires a tree search over possible token assignments. While this type of tree search is possible, it would take up significant runtime due to the large number of LLM forward passes required to try all mismatch combinations.
To simplify the procedure, we opt instead for a simpler, semi-greedy search that starts from the target model response and iteratively tries to replace mismatching target model tokens with their draft counterparts. If replacing a token affects the final answer, we consider this token important and keep the original (target model) version. If, however, replacing the token results in an equivalent answer, we deem this token unimportant, replace it with the draft model version and continue the search from the new sequence, with a different suffix and possibly a different $\mathcal{I}$. That way, we guarantee that the search algorithm is aligned with what happens during inference: the important tokens are generated with the target model and the unimportant ones are kept from the draft model. We summarize the resulting search procedure in Algorithm 1 and discuss some of its implications in Appendix A.
3.2 Classifier Training
@@
To find the total number of animal legs, we need to calculate the legs[$_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}\texttt{[total]}}}$ of each animal and then add them up.
- 2 dogs have 4[$_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}[2]}}$ legs each, so 2 dogs have 2 $_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}\texttt{[times]}}}$ 4 = 8 legs.
@@
Now[$_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}\texttt{[Adding]}}}$, let’s add the legs together[$_{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}\texttt{[of]}}}$ , we get 8 (from the dogs) + 2 (from the chicken) = 10 legs.
@@
# -$_{[\texttt{YOUR}]}$ coding: utf-8 -*-
[…]
@@
[…]
Once we gather a dataset of task-specific important tokens with Alg. 1, we can train a classifier that would detect such tokens for use during inference. This classifier can, in principle, be any type of model, from a simple linear model or decision tree to a fine-tuned transformer layer. However, in our work, we default to training lightweight linear models with existing LLM hidden states as features, since those would introduce the least overhead during inference. There are several important design choices that can affect the effectiveness of such a classifier: we address each one separately.
1. Which token representations to use: the hidden states that predicted the mismatched token, or the next hidden states that encode the mismatched token itself? In our experiments, we found that using the latter representations results in substantially greater classifier accuracy (see Appendix B). However, obtaining these representations comes with a caveat.
@@
2. Which token alternative to use? Since the classifier works best with the representations from encoding the mismatching token, it is natural to ask which token should be encoded: the draft token, the mismatching target token, or both? When analyzing this, we found that using both token representations comes with a slight increase in classifier accuracy (see Appendix B). However, obtaining these representations in practice would require running $\theta_{target}$ more than once during the verification stage, which would complicate inference and introduce performance overhead. For this reason, we opt to use only the draft token representations for the classifier, since those are already available in regular speculative decoding.
3. Which model provides feature representations? During the verification stage, we have access to both the draft and the target model hidden states: we can use either or both of them as the input. In practice, we found that concatenated draft and target model representations give slightly better results than just those of the target model, and using draft model representations alone is substantially worse. Since both representations are already available during inference, we opt to use each of them.
Classifier model & training. In this work, we train a simple logistic regression to detect important tokens. While a more complex model could achieve greater accuracy, logistic regression is significantly easier to deploy, has less runtime & memory overhead and needs less training data. Furthermore, it can be fused with the existing “LM head” layer of the draft and target LLMs, which would make its computation virtually free. To control overfitting, we perform a simple grid search over the $L_{2}$ regularization coefficient (“CC”) with a logarithmic grid. We report additional details in Appendix B.